If x

 If x

Grade:11

1 Answers

Kevin Nash
9 years ago
Given x < 0, y < 0
X + y + x/y = 1/2 and (x + y). x/y = -1/2
Let x + y = a and x/y = b . . . . . . . . . . (1)
∴ We get a + b = 1/2 and ab = -1/2
Solving these two, we get a + (1/2a) = 1/2
⇒ 2a2 – a – 1 = 0 ⇒ a = 1, -1/2
⇒ b = -1/2, 1
∴ ⇒ x + y = 1 and x/y = - 1/2
Or x + y = -1/2 and x/y = 1 But x, y < 0
∴ x + y < 0 ⇒ x + y = -1/2 and x/y = 1
On solving, we get x = -1/4 and y = -1/4.